the formula for the area of a parallelogram is A=bh where b represents base and h represents height. If the base of a parallelogram is 7feet more than the height, and the area is 60 square feet, then what are the measurements of the base and height?

thank you very much

Answer by scott8148(6628)

x(x + 7) = 60

x^2 + 7x - 60 = 0

(x + 12)(x - 5) = 0